OpenStudy (anonymous):

simplify and show work. -48/(7-9)^5 -2[1-5(2 - 6 + 3^2]

OpenStudy (anonymous):

-48/(-2)^5-2(1-5(5))=-48/-32-2(-24)=3/2+48=49.5